OpenStudy (anonymous):

Write the equation of the line perpendicular to the line x - 3y = 9 and passing through the point (3,5). A) y = -3x + 14 B. y = 2x C. y = -5x + 15 D.y = -4x + 20 E. y = -2x - 1 these are the options given

OpenStudy (anonymous):

so x - 3y = 9, so 3y = x+9 Y = 1/3 X + 3. can u tell slope of line now?

OpenStudy (primeralph):